Your firm currently has $ 52 million in debt outstanding with a 6 % interest rate. The terms of the loan require it to repay $ 13 million of the balance each year. Suppose the marginal corporate tax rate is 30 %​, and that the interest tax shields have the same risk as the loan. What is the present value of the interest tax shields from this​ debt?
Answer #1

Year 1
Beginning Balance=52
Interest=52*6%
Principal payment=13
Ending Balance=39

Year 2
Beginning Balance=39
Interest=39*6%
Principal payment=13
Ending Balance=26

Year 3
Beginning Balance=26
Interest=26*6%
Principal payment=13
Ending Balance=13

Year 4
Beginning Balance=13
Interest=13*6%
Principal payment=13
Ending Balance=0

Interest tax shields=6%*30%*(52/1.06+39/1.06^2+26/1.06^3+13/1.06^4)=2.08609 million=2086088.11047